Selling price $130 (same information as on page 6) Units in beginning inventory Units produced Units sold Units in ending inventory 6,100 6,000 100 Variable costs per unit: Direct materials Direct labor Variable manufacturing overhead $25 Variable selling and administrative S10 S33 530 Fixed costs: Fixed manufacturing overhead$91,500 Fixed selling and administrative $98,000 f. What is the net operating income for the month under absorption costing? i. Reconcile and explain any difference in the net income under the two different methods.
Answer #1

f

Income statement under absorption costing

Particulars Per unit Number of units Amount
Sales $       130 6000 $ 780,000
Minus:
Cost of goods sold
Direct materials $         33 6000 $ 198,000
Direct labor $         30 6000 $ 180,000
Variable manufacturing overhead $         25 6000 $ 150,000
Fixed manufacturing overhead $         15 6000 $   90,000
Total cost of goods sold $       103 $ 618,000
Gross profit $         27 $ 162,000
Minus:
Selling and administrative expenses:
Variable selling and administrative $         10 6000 $   60,000
Fixed selling and administrative $   98,000
Total selling and administrative expenses $ 158,000
Net operating income $      4,000

Fixed manufacturing overhead:

Under absorption costing fixed manufacturing overhead is allocated to the total production and then expensed to the portion of goods sold and the balance is capitalised as part of inventory.

Per unit fixed manufacturing overhead = total manufacturing overhead/ production = 91,500/6,100 = $15 per unit

Fixed manufacturing overhead expensed = units sold * per unit fixed manufacturing overhead = 15 * 6000= 90,000

Fixed manufacturing overhead forming part of closing inventory = closing inventory units * per unit fixed manufacturing overhead = 100 * 15 = 1500

Fixed Selling and administrative expenses:

Fixed selling and administrative expenses are considered period costs and expenses in the period incurred.

i

Reconciliation of net income under absorption costing and variable costing

Particulars Amount
Net operating income under absorption costing $       4,000
Less
Fixed manufacturing overhead capitalised as part of closing inventory under absorption costing $       1,500
Net operating income under variable costing $       2,500

Workings:

Income statement under variable costing:

Particulars Per unit Number of units Amount
Sales $       130 6000 $ 780,000
Minus:
Variable costs
Direct materials $         33 6000 $ 198,000
Direct labor $         30 6000 $ 180,000
Variable manufacturing overhead $         25 6000 $ 150,000
Variable selling and administrative expenses $         10 6000 $   60,000
Total variable costs $         98 $ 588,000
Contribution $         32 $ 192,000
Minus:
Fixed costs
Fixed manufacturing overhead $   91,500
Fixed selling and administrative $   98,000
Total fixed costs $ 189,500
Net operating income $      2,500

Note:

Fixed manufacturing costs:

Fixed manufacturing costs under variable costing are considered period costs and expensed in the period incurred under variable costing.

Reconciliation:

The difference in net income under both methods is due to capitalization of fixed manufacturing cost under absorption costing due to which net operating income is increased under absorption costing to that extent compared to variable costing.